OpenStudy (narissa):

simplify 5+6*4-10 divided by 5

OpenStudy (anonymous):

according to order of operations multiplication/division comes before addition/subtraction

OpenStudy (anonymous):

and if there are both multiplication and division in the equation, then you solve from left to right first...the same goes for addition/subtraction

OpenStudy (narissa):

so is it 27

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

nice

OpenStudy (narissa):

6*4=24-2 =5 or no?

OpenStudy (anonymous):

Whats with the star.

OpenStudy (narissa):

multiply

OpenStudy (anonymous):

it's the correct answer

OpenStudy (anonymous):

well a x is easier

OpenStudy (narissa):

thanks